The number of guppies in an aquarium is modelled by the function

K N(t) 5 12(1 1 0.04)t, where t is measured in weeks.

a) Describe what each part of the equation represents.

b) Determine the number of guppies in the aquarium after 10 weeks.

c) Will this equation always model the population in the aquarium?


1
Expert's answer
2021-02-02T03:58:17-0500

A) N(t) is the value function, the other part of equation is the function

B) N(10) = 5*(12*11*0.04t) = 5*(12*11*0.04 *10) = 264

C) it is a linear equation, thus, yes.